OpenStudy (anonymous):

Factor out the thre greatest common factor from the expression 6x^3yz^4 – 3x^2y^3z^4 – 18x^4y^2z^3.

OpenStudy (anonymous):

i kno the answer but here: purple flies down in scotland

OpenStudy (cwrw238):

gcf of numbers 6 3 and 18 = 3 terms in x: gcf = x^2 terms in y: gcf = y terms in z: gcf = z^3 GCF = 3x^2yz^3

OpenStudy (anonymous):

What is the most mistake-free way of determining the GCF you cam up with?